M<SXT=(4x+1), m<QXS=(2x-2), and m<QXT=125. Find the value of x and m<QXS
1
Expert's answer
2020-09-22T16:06:36-0400

"m<\textrm{SXT} = m<\textrm{QXS} + m<\textrm{QXT}\\



\Rightarrow 125 + 2x - 2 = 4x + 1 \\

125 - 2 - 1 = 4x - 2x \\

2x = 122\\

x = 61\degree\\



\therefore m<\textrm{QXS} = 2x - 2 = 2(61) - 2 = 122 - 2 = 120\degree"